OpenStudy (anonymous):

You found a bag of dollar bills and you return it to the police. Later on, an officer tells you that the bag contained 3 kinds of bills an O amount of $1 bills, F amount of $5, and L amount of $50 dollar bills. The bag had 110 bills worth $1620. Moreover, there were 3 times as many $50 bills than $5 bills. Write a system of 3 linear equations. .. I'm pretty sure the first two are O+F+L=110, and O+5F+50L=1620, but I can't figure out the last one. I need to put the equations in a matrix and I've tried everything to have it make sense and it doesn't :S

OpenStudy (anonymous):

O + F + L = 110 O + 5F + 50L = 1620 L = 3F

OpenStudy (anonymous):

Three times as many 50 bills as 5 bills: L = 3*F